# Harmonizing Generalization and Specialization: Uncertainty-Informed Collaborative Learning for Semi-supervised Medical Image Segmentation

### v1

@@ 4.1 Experimental Settings
Left Thalamus (LT) | Left Hippocampus (LH) | Right Thalamus (RT) | Right Hippocampus (RH) | Average | |||||||||||||||||
| DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | ||||
| SAM(ICCV’23) [5] |
39.604.47 | 24.783.41 | 77.829.59 | 9.790.53 | 20.842.91 | 11.661.81 | 95.541.10 | 11.030.73 | 51.912.53 | 35.092.26 | 15.032.70 | 8.330.55 | 39.953.73 | 25.032.85 | 15.161.28 | 10.000.67 | 38.0711.66 | 24.148.73 | 50.8936.69 | 9.791.15 | ||
| MedSAM(NC’24) [6] | 78.417.30 | 65.099.97 | 7.512.18 | 2.060.88 | 75.578.90 | 61.5411.00 | 6.851.99 | 2.220.80 | 75.099.10 | 60.829.65 | 8.132.10 | 2.490.93 | 76.888.97 | 63.2811.27 | 6.711.87 | 2.130.79 | 76.498.78 | 62.6810.71 | 7.302.13 | 2.230.87 | |||
| UNet(MICCAI’15) [31] |
83.549.77 | 75.2411.65 | 14.1010.55 | 4.533.18 | 47.6524.90 | 35.1821.50 | 64.6923.62 | 29.6715.97 | 77.8912.76 | 70.6813.85 | 8.006.29 | 3.322.77 | 16.3422.34 | 12.0617.96 | 17.2223.84 | 7.6314.13 | 56.3636.25 | 48.2934.05 | 26.0030.68 | 11.2916.34 | ||
| nnU-Net(Nat. Methods’21) [32] | 93.35 2.54 | 87.64 4.27 | 5.93 10.94 | 2.10 3.04 | 87.11 7.80 | 77.96 10.86 | 14.08 26.28 | 4.11 6.96 | 93.27 2.70 | 87.52 4.50 | 5.45 9.95 | 1.79 2.61 | 86.63 7.37 | 77.19 9.97 | 41.17 38.65 | 13.19 12.28 | 90.09 6.59 | 82.58 9.49 | 16.66 29.08 | 5.30 8.77 | |||
| MT(NeurIPS’17) [3] | 81.1019.39 | 72.8021.59 | 20.8420.87 | 7.177.38 | 71.7720.96 | 60.2022.34 | 43.2130.77 | 17.2413.07 | 80.4622.78 | 72.9323.88 | 15.5017.30 | 4.805.57 | 57.1231.59 | 49.4329.85 | 19.7127.91 | 7.0413.70 | 72.6127.18 | 63.8427.45 | 24.8229.77 | 9.0612.62 | |||
| UAMT(MICCAI’19) [4] | 83.4912.92 | 74.0015.12 | 19.2014.85 | 6.324.92 | 67.0122.58 | 55.2423.07 | 39.6830.69 | 15.7613.17 | 79.1919.36 | 70.6420.28 | 12.8513.38 | 3.774.37 | 55.3830.49 | 46.0528.71 | 28.4529.15 | 11.0314.42 | 71.2726.61 | 61.4826.50 | 25.0528.65 | 9.2212.50 | |||
| URPC(MIA’22) [15] | 70.4025.24 | 62.9325.69 | 11.1512.15 | 4.354.77 | 61.9226.94 | 51.4826.28 | 37.4428.94 | 15.9215.04 | 66.0727.17 | 59.4626.88 | 10.0810.26 | 4.135.63 | 54.2330.75 | 45.5428.94 | 26.4930.73 | 12.1516.28 | 63.1530.93 | 54.8630.16 | 21.2927.53 | 9.1413.63 | |||
| BCP(CVPR’23) [13] | 92.064.98 | 85.806.97 | 3.476.31 | 1.171.40 | 85.498.71 | 75.8211.26 | 7.1410.98 | 2.393.93 | 91.135.89 | 84.288.12 | 3.234.83 | 1.201.47 | 88.2410.00 | 80.3911.54 | 7.8912.30 | 2.634.70 | 89.238.93 | 81.5711.28 | 5.4310.94 | 1.853.74 | |||
| ABD(CVPR’24) [12] | 94.183.14 | 89.303.30 | 2.863.60 | 0.941.11 | 89.491.34 | 81.351.38 | 5.708.23 | 1.791.92 | 94.382.84 | 89.673.17 | 3.123.53 | 0.990.99 | 90.010.80 | 82.141.08 | 5.005.83 | 1.541.91 | 92.011.56 | 85.611.62 | 4.174.72 | 1.311.10 | |||
| CPC-SAM(MICCAI’24) [16] | 93.831.57 | 88.422.76 | 2.130.51 | 0.790.25 | 88.392.64 | 79.304.21 | 2.890.95 | 1.010.36 | 93.871.62 | 88.492.87 | 2.251.34 | 0.850.42 | 89.223.70 | 80.735.81 | 2.621.02 | 0.900.26 | 91.333.59 | 84.245.91 | 2.471.04 | 0.890.34 | |||
| UnCoL(ours) | 95.541.31 | 91.502.31 | 1.521.93 | 0.540.39 | 92.363.28 | 85.975.19 | 2.904.06 | 0.811.22 | 95.461.36 | 91.352.42 | 1.470.55 | 0.540.15 | 92.903.18 | 86.925.05 | 2.825.06 | 0.781.25 | 94.072.92 | 88.934.78 | 2.184.63 | 0.671.28 | |||
| UNet(MICCAI’15) [31] |
86.338.15 | 77.2510.37 | 10.029.08 | 3.002.44 | 48.9627.54 | 38.1024.79 | 53.9528.02 | 25.5717.31 | 86.527.24 | 77.439.72 | 10.397.79 | 3.312.13 | 43.7428.93 | 34.0725.47 | 21.6620.83 | 7.479.57 | 66.3930.16 | 56.7129.29 | 24.0128.74 | 9.8415.09 | ||
| nnU-Net(Nat. Methods’21) [32] | 92.71 4.65 | 86.84 7.08 | 6.57 10.65 | 2.11 3.01 | 88.84 7.25 | 80.63 9.44 | 7.59 15.96 | 2.31 4.49 | 92.03 6.23 | 86.06 8.59 | 3.91 5.76 | 1.37 1.53 | 88.91 5.90 | 80.51 8.73 | 31.18 34.45 | 9.45 10.13 | 90.62 6.69 | 83.51 9.43 | 12.31 22.94 | 3.81 6.68 | |||
| MT(NeurIPS’17) [3] | 90.2310.05 | 84.0012.85 | 6.659.56 | 2.092.98 | 79.7720.48 | 70.6222.58 | 21.5929.31 | 7.9211.70 | 89.1714.84 | 82.9917.03 | 6.7110.62 | 2.033.25 | 76.9224.65 | 68.0325.64 | 16.4924.93 | 5.439.29 | 84.0220.28 | 76.4122.09 | 12.8622.86 | 4.378.88 | |||
| UAMT(MICCAI’19) [4] | 89.2613.78 | 82.7217.07 | 8.3211.67 | 2.563.72 | 78.9323.31 | 70.3824.81 | 22.0231.79 | 8.1113.04 | 88.2716.89 | 81.9919.33 | 7.9711.55 | 2.363.47 | 76.2128.95 | 68.5429.12 | 16.1325.21 | 5.4910.19 | 83.1622.76 | 75.9124.28 | 13.6123.20 | 4.639.27 | |||
| URPC(MIA’22) [15] | 88.2815.15 | 82.1416.76 | 4.008.34 | 0.972.23 | 80.4722.35 | 71.9123.71 | 14.6125.01 | 4.949.69 | 87.9715.07 | 82.0916.54 | 3.808.26 | 0.842.31 | 78.2824.91 | 70.6625.48 | 11.8324.52 | 4.0410.09 | 83.7521.51 | 76.7022.59 | 8.5619.99 | 2.707.76 | |||
| BCP(CVPR’23) [13] | 93.072.51 | 87.164.18 | 2.271.80 | 0.860.40 | 90.175.09 | 82.587.25 | 4.638.27 | 1.462.79 | 92.393.75 | 86.125.87 | 2.351.19 | 0.930.43 | 90.666.59 | 83.708.62 | 5.8414.17 | 1.643.59 | 91.575.26 | 84.897.41 | 3.778.91 | 1.222.48 | |||
| ABD(CVPR’24) [12] | 95.280.83 | 91.081.20 | 2.131.80 | 0.680.63 | 90.911.07 | 83.671.19 | 3.294.37 | 1.081.82 | 95.331.47 | 91.231.84 | 1.881.36 | 0.620.43 | 91.073.05 | 84.192.32 | 3.326.06 | 1.072.54 | 93.151.44 | 87.541.26 | 2.652.81 | 0.861.01 | |||
| CPC-SAM(MICCAI’24) [16] | 95.251.27 | 90.952.28 | 1.900.60 | 0.580.17 | 89.113.74 | 80.565.76 | 2.660.67 | 0.970.77 | 95.051.46 | 90.612.63 | 1.920.59 | 0.590.18 | 90.192.68 | 82.234.33 | 2.540.66 | 0.820.19 | 92.403.73 | 86.096.20 | 2.250.72 | 0.740.45 | |||
| UnCoL(ours) | 95.631.13 | 91.642.02 | 2.074.63 | 0.661.03 | 93.012.34 | 87.033.94 | 2.132.17 | 0.570.45 | 95.731.07 | 91.831.93 | 1.370.45 | 0.500.14 | 93.322.12 | 87.563.63 | 2.694.62 | 0.730.89 | 94.422.17 | 89.513.77 | 2.064.07 | 0.610.87 | |||
| UNet(MICCAI’15) [31] |
94.348.80 | 90.389.79 | 2.805.99 | 0.952.16 | 88.6916.05 | 82.2317.16 | 4.5911.61 | 1.313.94 | 94.716.91 | 90.698.50 | 2.615.65 | 0.911.95 | 91.0510.24 | 84.8212.38 | 11.7924.31 | 3.878.62 | 92.2011.90 | 87.0313.41 | 5.4515.14 | 1.765.32 | ||
| nnU-Net(Nat. Methods’21) [32] | 95.05 7.56 | 91.30 7.86 | 1.88 3.77 | 0.71 1.83 | 92.65 6.55 | 86.87 7.73 | 2.96 7.47 | 1.10 4.08 | 95.09 7.07 | 91.30 7.54 | 1.80 3.28 | 0.69 1.54 | 93.07 6.38 | 87.57 7.78 | 3.07 8.06 | 1.17 4.30 | 93.96 7.13 | 89.26 8.12 | 2.43 6.15 | 0.92 3.27 | |||
| Method |
Metrics | |||||
| DSC | Jaccard | 95HD | ASD | ||||
| SAM(ICCV’23) [5] |
18.137.37 | 10.164.50 | 17.866.57 | 4.271.80 | ||
| MedSAM(NC’24) [6] | 70.924.08 | 55.104.82 | 7.711.59 | 2.680.51 | |||
| SAM-Med3D [7] | 79.374.55 | 66.026.22 | 5.521.79 | 1.550.54 | |||
| VNet(3DV’16) [33] |
56.9616.50 | 41.3813.66 | 18.6912.60 | 5.383.90 | ||
| nnU-Net(Nat. Methods’21) [32] | 49.06 20.36 | 34.93 16.73 | 23.66 10.13 | 8.62 4.87 | |||
| Swin-UNETR(MICCAI’21) [2] | 67.1110.59 | 51.4211.51 | 12.3511.99 | 2.913.10 | |||
| MT(NeurIPS’17) [3] | 64.2811.62 | 48.4012.10 | 28.1914.54 | 8.925.15 | |||
| UA-MT(MICCAI’19) [4] | 62.048.07 | 45.468.44 | 33.5415.97 | 11.265.22 | |||
| MC-Net+(MIA’22) [22] | 64.2613.61 | 48.8214.82 | 33.6323.19 | 13.029.57 | |||
| FUSSNet(MICCAI’22) [20] | 59.3212.50 | 43.2412.25 | 30.8313.64 | 10.295.94 | |||
| CauSSL(ICCV’23) [19] | 64.3513.30 | 48.7513.54 | 23.0111.60 | 7.774.13 | |||
| AC-MT(MIA’23) [11] | 66.508.72 | 50.439.42 | 21.9611.48 | 7.313.90 | |||
| BCP(CVPR’23) [13] | 80.738.29 | 68.4810.17 | 7.677.35 | 2.382.41 | |||
| Su et al.(MIA’24) [14] | 71.299.70 | 56.2511.50 | 18.1912.57 | 5.974.45 | |||
| SemiSAM+(arXiv’25) [23] | 64.7210.33 | 48.6510.52 | 20.5310.39 | 6.103.19 | |||
| UnCoL(ours) | 80.405.82 | 67.607.83 | 6.863.96 | 1.850.94 | |||
| VNet(3DV’16) [33] |
70.5211.40 | 55.6213.12 | 10.436.05 | 2.281.83 | ||
| nnU-Net(Nat. Methods’21) [32] | 60.06 17.06 | 45.08 15.95 | 16.65 9.14 | 5.51 4.04 | |||
| Swin-UNETR(MICCAI’21) [2] | 75.427.73 | 61.149.62 | 5.123.52 | 1.000.98 | |||
| MT(NeurIPS’17) [3] | 74.978.21 | 60.6310.17 | 12.188.72 | 3.592.56 | |||
| UA-MT(MICCAI’19) [4] | 76.457.97 | 62.5310.12 | 9.475.25 | 2.991.54 | |||
| MC-Net+(MIA’22) [22] | 77.988.76 | 64.6910.95 | 8.976.50 | 2.631.75 | |||
| FUSSNet(MICCAI’22) [20] | 77.268.15 | 63.6410.56 | 12.669.57 | 4.023.09 | |||
| CauSSL(ICCV’23) [19] | 77.449.27 | 64.0211.17 | 11.9610.49 | 3.523.10 | |||
| AC-MT(MIA’23) [11] | 76.398.63 | 62.5510.78 | 10.807.04 | 3.302.23 | |||
| BCP(CVPR’23) [13] | 82.935.56 | 71.217.65 | 6.305.46 | 1.831.34 | |||
| Su et al.(MIA’24) [14] | 81.016.04 | 68.508.26 | 6.323.28 | 1.721.14 | |||
| SemiSAM+(arXiv’25) [23] | 75.976.96 | 61.768.97 | 8.895.76 | 2.571.37 | |||
| UnCoL(ours) | 83.164.80 | 71.466.79 | 5.443.26 | 1.440.90 | |||
| VNet(3DV’16) [33] |
81.896.33 | 69.788.46 | 9.338.26 | 2.391.99 | ||
| nnU-Net(Nat. Methods’21) [32] | 79.93 7.05 | 67.91 8.72 | 6.32 3.35 | 1.94 0.89 | |||
| Swin-UNETR(MICCAI’21) [2] | 83.493.98 | 71.865.75 | 2.351.91 | 0.450.32 | |||
| SegVol(NeurIPS’24) [34] | 58.8316.46 | 42.2316.54 | 21.6219.64 | 5.434.85 | |||
| Methods |
True Lumen (TL) | False Lumen (FL) | Average | |||||||||||
| DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | ||||
| SAM(ICCV’23) [5] |
19.184.48 | 10.992.89 | 16.147.61 | 5.641.30 | 18.173.53 | 10.312.25 | 16.423.98 | 5.551.03 | 18.684.07 | 10.652.62 | 16.286.15 | 5.601.18 | ||
| MedSAM(NC’24) [6] | 47.8811.68 | 32.239.70 | 7.892.75 | 2.890.95 | 40.356.21 | 25.484.91 | 11.732.43 | 4.540.81 | 44.1210.13 | 28.868.43 | 9.813.26 | 3.711.22 | |||
| SAM-Med3D [7] | 33.5620.60 | 22.2315.73 | 13.556.71 | 4.422.52 | 25.6616.18 | 15.8611.85 | 12.913.93 | 4.471.06 | 29.6119.11 | 19.0414.43 | 13.235.58 | 4.441.94 | |||
| SegVol(NeurIPS’24) [34] | 34.4115.24 | 21.8511.73 | 18.687.39 | 5.902.93 | 30.8716.84 | 19.2212.29 | 19.6410.90 | 6.013.37 | 32.6416.15 | 20.5312.09 | 19.169.32 | 5.963.16 | |||
| VNet(3DV’16) [33] |
63.2514.37 | 48.4114.05 | 13.3410.49 | 3.204.00 | 48.4523.80 | 36.4120.80 | 15.0412.32 | 4.095.00 | 55.8521.20 | 42.4118.96 | 14.1911.74 | 3.654.56 | ||
| nnU-Net(Nat. Methods’21) [32] | 73.14 11.63 | 59.04 13.44 | 6.25 3.96 | 1.67 1.03 | 64.07 20.09 | 50.34 20.28 | 8.10 9.53 | 2.49 2.94 | 68.61 17.28 | 54.69 18.04 | 7.17 8.04 | 2.08 2.26 | |||
| Swin-UNETR(MICCAI’21) [2] | 64.74 9.35 | 48.74 9.81 | 6.02 5.25 | 1.75 2.22 | 59.74 13.70 | 44.04 13.64 | 5.85 6.51 | 1.57 1.72 | 62.24 12.28 | 46.39 12.44 | 5.94 6.33 | 1.66 2.16 | |||
| MT(NeurIPS’17) [3] | 70.419.54 | 55.4010.91 | 5.894.75 | 1.460.79 | 58.9218.60 | 44.6817.82 | 9.074.57 | 2.541.58 | 64.6715.93 | 50.0415.83 | 7.486.06 | 2.001.43 | |||
[…]

### v2

@@ 4.1 Experimental Settings
| Method | Labeled Ratio | Left Thalamus (LT) | Left Hippocampus (LH) | Right Thalamus (RT) | Right Hippocampus (RH) | Average | |||||||||||||||
| DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | ||
| SAM(ICCV’23) [18] | 0/289 () | 39.604.47 | 24.783.41 | 77.829.59 | 9.790.53 | 20.842.91 | 11.661.81 | 95.541.10 | 11.030.73 | 51.912.53 | 35.092.26 | 15.032.70 | 8.330.55 | 39.953.73 | 25.032.85 | 15.161.28 | 10.000.67 | 38.0711.66 | 24.148.73 | 50.8936.69 | 9.791.15 |
| MedSAM(NC’24) [23] | 78.417.30 | 65.099.97 | 7.512.18 | 2.060.88 | 75.578.90 | 61.5411.00 | 6.851.99 | 2.220.80 | 75.099.10 | 60.829.65 | 8.132.10 | 2.490.93 | 76.888.97 | 63.2811.27 | 6.711.87 | 2.130.79 | 76.498.78 | 62.6810.71 | 7.302.13 | 2.230.87 | |
| UNet(MICCAI’15) [30] | 15/289 () | 83.549.77 | 75.2411.65 | 14.1010.55 | 4.533.18 | 47.6524.90 | 35.1821.50 | 64.6923.62 | 29.6715.97 | 77.8912.76 | 70.6813.85 | 8.006.29 | 3.322.77 | 16.3422.34 | 12.0617.96 | 17.2223.84 | 7.6314.13 | 56.3636.25 | 48.2934.05 | 26.0030.68 | 11.2916.34 |
| nnU-Net(Nat. Methods’21) [17] | 93.35 2.54 | 87.64 4.27 | 5.93 10.94 | 2.10 3.04 | 87.11 7.80 | 77.96 10.86 | 14.08 26.28 | 4.11 6.96 | 93.27 2.70 | 87.52 4.50 | 5.45 9.95 | 1.79 2.61 | 86.63 7.37 | 77.19 9.97 | 41.17 38.65 | 13.19 12.28 | 90.09 6.59 | 82.58 9.49 | 16.66 29.08 | 5.30 8.77 | |
| MT(NeurIPS’17) [33] | 81.1019.39 | 72.8021.59 | 20.8420.87 | 7.177.38 | 71.7720.96 | 60.2022.34 | 43.2130.77 | 17.2413.07 | 80.4622.78 | 72.9323.88 | 15.5017.30 | 4.805.57 | 57.1231.59 | 49.4329.85 | 19.7127.91 | 7.0413.70 | 72.6127.18 | 63.8427.45 | 24.8229.77 | 9.0612.62 | |
| UAMT(MICCAI’19) [41] | 83.4912.92 | 74.0015.12 | 19.2014.85 | 6.324.92 | 67.0122.58 | 55.2423.07 | 39.6830.69 | 15.7613.17 | 79.1919.36 | 70.6420.28 | 12.8513.38 | 3.774.37 | 55.3830.49 | 46.0528.71 | 28.4529.15 | 11.0314.42 | 71.2726.61 | 61.4826.50 | 25.0528.65 | 9.2212.50 | |
| URPC(MIA’22) [22] | 70.4025.24 | 62.9325.69 | 11.1512.15 | 4.354.77 | 61.9226.94 | 51.4826.28 | 37.4428.94 | 15.9215.04 | 66.0727.17 | 59.4626.88 | 10.0810.26 | 4.135.63 | 54.2330.75 | 45.5428.94 | 26.4930.73 | 12.1516.28 | 63.1530.93 | 54.8630.16 | 21.2927.53 | 9.1413.63 | |
| BCP(CVPR’23) [3] | 92.064.98 | 85.806.97 | 3.476.31 | 1.171.40 | 85.498.71 | 75.8211.26 | 7.1410.98 | 2.393.93 | 91.135.89 | 84.288.12 | 3.234.83 | 1.201.47 | 88.2410.00 | 80.3911.54 | 7.8912.30 | 2.634.70 | 89.238.93 | 81.5711.28 | 5.4310.94 | 1.853.74 | |
| ABD(CVPR’24) [7] | 94.183.14 | 89.303.30 | 2.863.60 | 0.941.11 | 89.491.34 | 81.351.38 | 5.708.23 | 1.791.92 | 94.382.84 | 89.673.17 | 3.123.53 | 0.990.99 | 90.010.80 | 82.141.08 | 5.005.83 | 1.541.91 | 92.011.56 | 85.611.62 | 4.174.72 | 1.311.10 | |
| CPC-SAM(MICCAI’24) [28] | 93.831.57 | 88.422.76 | 2.130.51 | 0.790.25 | 88.392.64 | 79.304.21 | 2.890.95 | 1.010.36 | 93.871.62 | 88.492.87 | 2.251.34 | 0.850.42 | 89.223.70 | 80.735.81 | 2.621.02 | 0.900.26 | 91.333.59 | 84.245.91 | 2.471.04 | 0.890.34 | |
| UnCoL(ours) | 95.541.31 | 91.502.31 | 1.521.93 | 0.540.39 | 92.363.28 | 85.975.19 | 2.904.06 | 0.811.22 | 95.461.36 | 91.352.42 | 1.470.55 | 0.540.15 | 92.903.18 | 86.925.05 | 2.825.06 | 0.781.25 | 94.072.92 | 88.934.78 | 2.184.63 | 0.671.28 | |
| UNet(MICCAI’15) [30] | 29/289 () | 86.338.15 | 77.2510.37 | 10.029.08 | 3.002.44 | 48.9627.54 | 38.1024.79 | 53.9528.02 | 25.5717.31 | 86.527.24 | 77.439.72 | 10.397.79 | 3.312.13 | 43.7428.93 | 34.0725.47 | 21.6620.83 | 7.479.57 | 66.3930.16 | 56.7129.29 | 24.0128.74 | 9.8415.09 |
| nnU-Net(Nat. Methods’21) [17] | 92.71 4.65 | 86.84 7.08 | 6.57 10.65 | 2.11 3.01 | 88.84 7.25 | 80.63 9.44 | 7.59 15.96 | 2.31 4.49 | 92.03 6.23 | 86.06 8.59 | 3.91 5.76 | 1.37 1.53 | 88.91 5.90 | 80.51 8.73 | 31.18 34.45 | 9.45 10.13 | 90.62 6.69 | 83.51 9.43 | 12.31 22.94 | 3.81 6.68 | |
| MT(NeurIPS’17) [33] | 90.2310.05 | 84.0012.85 | 6.659.56 | 2.092.98 | 79.7720.48 | 70.6222.58 | 21.5929.31 | 7.9211.70 | 89.1714.84 | 82.9917.03 | 6.7110.62 | 2.033.25 | 76.9224.65 | 68.0325.64 | 16.4924.93 | 5.439.29 | 84.0220.28 | 76.4122.09 | 12.8622.86 | 4.378.88 | |
| UAMT(MICCAI’19) [41] | 89.2613.78 | 82.7217.07 | 8.3211.67 | 2.563.72 | 78.9323.31 | 70.3824.81 | 22.0231.79 | 8.1113.04 | 88.2716.89 | 81.9919.33 | 7.9711.55 | 2.363.47 | 76.2128.95 | 68.5429.12 | 16.1325.21 | 5.4910.19 | 83.1622.76 | 75.9124.28 | 13.6123.20 | 4.639.27 | |
| URPC(MIA’22) [22] | 88.2815.15 | 82.1416.76 | 4.008.34 | 0.972.23 | 80.4722.35 | 71.9123.71 | 14.6125.01 | 4.949.69 | 87.9715.07 | 82.0916.54 | 3.808.26 | 0.842.31 | 78.2824.91 | 70.6625.48 | 11.8324.52 | 4.0410.09 | 83.7521.51 | 76.7022.59 | 8.5619.99 | 2.707.76 | |
| BCP(CVPR’23) [3] | 93.072.51 | 87.164.18 | 2.271.80 | 0.860.40 | 90.175.09 | 82.587.25 | 4.638.27 | 1.462.79 | 92.393.75 | 86.125.87 | 2.351.19 | 0.930.43 | 90.666.59 | 83.708.62 | 5.8414.17 | 1.643.59 | 91.575.26 | 84.897.41 | 3.778.91 | 1.222.48 | |
| ABD(CVPR’24) [7] | 95.280.83 | 91.081.20 | 2.131.80 | 0.680.63 | 90.911.07 | 83.671.19 | 3.294.37 | 1.081.82 | 95.331.47 | 91.231.84 | 1.881.36 | 0.620.43 | 91.073.05 | 84.192.32 | 3.326.06 | 1.072.54 | 93.151.44 | 87.541.26 | 2.652.81 | 0.861.01 | |
| CPC-SAM(MICCAI’24) [28] | 95.251.27 | 90.952.28 | 1.900.60 | 0.580.17 | 89.113.74 | 80.565.76 | 2.660.67 | 0.970.77 | 95.051.46 | 90.612.63 | 1.920.59 | 0.590.18 | 90.192.68 | 82.234.33 | 2.540.66 | 0.820.19 | 92.403.73 | 86.096.20 | 2.250.72 | 0.740.45 | |
| UnCoL(ours) | 95.631.13 | 91.642.02 | 2.074.63 | 0.661.03 | 93.012.34 | 87.033.94 | 2.132.17 | 0.570.45 | 95.731.07 | 91.831.93 | 1.370.45 | 0.500.14 | 93.322.12 | 87.563.63 | 2.694.62 | 0.730.89 | 94.422.17 | 89.513.77 | 2.064.07 | 0.610.87 | |
| UNet(MICCAI’15) [30] | 289/289 () | 94.348.80 | 90.389.79 | 2.805.99 | 0.952.16 | 88.6916.05 | 82.2317.16 | 4.5911.61 | 1.313.94 | 94.716.91 | 90.698.50 | 2.615.65 | 0.911.95 | 91.0510.24 | 84.8212.38 | 11.7924.31 | 3.878.62 | 92.2011.90 | 87.0313.41 | 5.4515.14 | 1.765.32 |
| nnU-Net(Nat. Methods’21) [17] | 95.05 7.56 | 91.30 7.86 | 1.88 3.77 | 0.71 1.83 | 92.65 6.55 | 86.87 7.73 | 2.96 7.47 | 1.10 4.08 | 95.09 7.07 | 91.30 7.54 | 1.80 3.28 | 0.69 1.54 | 93.07 6.38 | 87.57 7.78 | 3.07 8.06 | 1.17 4.30 | 93.96 7.13 | 89.26 8.12 | 2.43 6.15 | 0.92 3.27 | |
| Method | Labeled Ratio | Metrics | |||
| DSC | Jaccard | 95HD | ASD | ||
| SAM(ICCV’23) [18] | 0/62 () | 18.137.37 | 10.164.50 | 17.866.57 | 4.271.80 |
| MedSAM(NC’24) [23] | 70.924.08 | 55.104.82 | 7.711.59 | 2.680.51 | |
| SAM-Med3D [34] | 79.374.55 | 66.026.22 | 5.521.79 | 1.550.54 | |
| VNet(3DV’16) [29] | 6/62 () | 56.9616.50 | 41.3813.66 | 18.6912.60 | 5.383.90 |
| nnU-Net(Nat. Methods’21) [17] | 49.06 20.36 | 34.93 16.73 | 23.66 10.13 | 8.62 4.87 | |
| Swin-UNETR(MICCAI’21) [14] | 67.1110.59 | 51.4211.51 | 12.3511.99 | 2.913.10 | |
| MT(NeurIPS’17) [33] | 64.2811.62 | 48.4012.10 | 28.1914.54 | 8.925.15 | |
| UA-MT(MICCAI’19) [41] | 62.048.07 | 45.468.44 | 33.5415.97 | 11.265.22 | |
| MC-Net+(MIA’22) [37] | 64.2613.61 | 48.8214.82 | 33.6323.19 | 13.029.57 | |
| FUSSNet(MICCAI’22) [38] | 59.3212.50 | 43.2412.25 | 30.8313.64 | 10.295.94 | |
| CauSSL(ICCV’23) [27] | 64.3513.30 | 48.7513.54 | 23.0111.60 | 7.774.13 | |
| AC-MT(MIA’23) [39] | 66.508.72 | 50.439.42 | 21.9611.48 | 7.313.90 | |
| BCP(CVPR’23) [3] | 80.738.29 | 68.4810.17 | 7.677.35 | 2.382.41 | |
| Su et al.(MIA’24) [32] | 71.299.70 | 56.2511.50 | 18.1912.57 | 5.974.45 | |
| SemiSAM+(arXiv’25) [44] | 64.7210.33 | 48.6510.52 | 20.5310.39 | 6.103.19 | |
| UnCoL(ours) | 80.405.82 | 67.607.83 | 6.863.96 | 1.850.94 | |
| VNet(3DV’16) [29] | 12/62 () | 70.5211.40 | 55.6213.12 | 10.436.05 | 2.281.83 |
| nnU-Net(Nat. Methods’21) [17] | 60.06 17.06 | 45.08 15.95 | 16.65 9.14 | 5.51 4.04 | |
| Swin-UNETR(MICCAI’21) [14] | 75.427.73 | 61.149.62 | 5.123.52 | 1.000.98 | |
| MT(NeurIPS’17) [33] | 74.978.21 | 60.6310.17 | 12.188.72 | 3.592.56 | |
| UA-MT(MICCAI’19) [41] | 76.457.97 | 62.5310.12 | 9.475.25 | 2.991.54 | |
| MC-Net+(MIA’22) [37] | 77.988.76 | 64.6910.95 | 8.976.50 | 2.631.75 | |
| FUSSNet(MICCAI’22) [38] | 77.268.15 | 63.6410.56 | 12.669.57 | 4.023.09 | |
| CauSSL(ICCV’23) [27] | 77.449.27 | 64.0211.17 | 11.9610.49 | 3.523.10 | |
| AC-MT(MIA’23) [39] | 76.398.63 | 62.5510.78 | 10.807.04 | 3.302.23 | |
| BCP(CVPR’23) [3] | 82.935.56 | 71.217.65 | 6.305.46 | 1.831.34 | |
| Su et al.(MIA’24) [32] | 81.016.04 | 68.508.26 | 6.323.28 | 1.721.14 | |
| SemiSAM+(arXiv’25) [44] | 75.976.96 | 61.768.97 | 8.895.76 | 2.571.37 | |
| UnCoL(ours) | 83.164.80 | 71.466.79 | 5.443.26 | 1.440.90 | |
| VNet(3DV’16) [29] | 62/62 () | 81.896.33 | 69.788.46 | 9.338.26 | 2.391.99 |
| nnU-Net(Nat. Methods’21) [17] | 79.93 7.05 | 67.91 8.72 | 6.32 3.35 | 1.94 0.89 | |
| Swin-UNETR(MICCAI’21) [14] | 83.493.98 | 71.865.75 | 2.351.91 | 0.450.32 | |
| SegVol(NeurIPS’24) [10] | 58.8316.46 | 42.2316.54 | 21.6219.64 | 5.434.85 | |
| Methods | Labeled Ratio | True Lumen (TL) | False Lumen (FL) | Average | |||||||||
| DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | ||
| SAM(ICCV’23) [18] | 0/70 () | 19.184.48 | 10.992.89 | 16.147.61 | 5.641.30 | 18.173.53 | 10.312.25 | 16.423.98 | 5.551.03 | 18.684.07 | 10.652.62 | 16.286.15 | 5.601.18 |
| MedSAM(NC’24) [23] | 47.8811.68 | 32.239.70 | 7.892.75 | 2.890.95 | 40.356.21 | 25.484.91 | 11.732.43 | 4.540.81 | 44.1210.13 | 28.868.43 | 9.813.26 | 3.711.22 | |
| SAM-Med3D [34] | 33.5620.60 | 22.2315.73 | 13.556.71 | 4.422.52 | 25.6616.18 | 15.8611.85 | 12.913.93 | 4.471.06 | 29.6119.11 | 19.0414.43 | 13.235.58 | 4.441.94 | |
| SegVol(NeurIPS’24) [10] | 34.4115.24 | 21.8511.73 | 18.687.39 | 5.902.93 | 30.8716.84 | 19.2212.29 | 19.6410.90 | 6.013.37 | 32.6416.15 | 20.5312.09 | 19.169.32 | 5.963.16 | |
| VNet(3DV’16) [29] | 7/70 () | 63.2514.37 | 48.4114.05 | 13.3410.49 | 3.204.00 | 48.4523.80 | 36.4120.80 | 15.0412.32 | 4.095.00 | 55.8521.20 | 42.4118.96 | 14.1911.74 | 3.654.56 |
| nnU-Net(Nat. Methods’21) [17] | 73.14 11.63 | 59.04 13.44 | 6.25 3.96 | 1.67 1.03 | 64.07 20.09 | 50.34 20.28 | 8.10 9.53 | 2.49 2.94 | 68.61 17.28 | 54.69 18.04 | 7.17 8.04 | 2.08 2.26 | |
| Swin-UNETR(MICCAI’21) [14] | 64.74 9.35 | 48.74 9.81 | 6.02 5.25 | 1.75 2.22 | 59.74 13.70 | 44.04 13.64 | 5.85 6.51 | 1.57 1.72 | 62.24 12.28 | 46.39 12.44 | 5.94 6.33 | 1.66 2.16 | |
| MT(NeurIPS’17) [33] | 70.419.54 | 55.4010.91 | 5.894.75 | 1.460.79 | 58.9218.60 | 44.6817.82 | 9.074.57 | 2.541.58 | 64.6715.93 | 50.0415.83 | 7.486.06 | 2.001.43 | |
[…]